archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Orange s webservices for schedule management The command to create new entry is impacted We will contact Orange to get the problem resolved Date Friday 11 July 2014 11 44AM Reason for closing Done Comments 1 Comment by OVH Friday

    Original URL path: http://status.ovh.es/?do=details&id=7262 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    2014 15 12PM Reason for closing Done Comments 2 Comment by OVH Thursday 10 July 2014 10 41AM Following a fibre disconnect on the optical loop in Paris signal is no longer reaching the TH2 site The optical loop is redundant but it seems that a loop compenent did not properly manage the disconnect We will investigate Only the FT collect is impacted Security is doing its job for subscribers

    Original URL path: http://status.ovh.es/?do=details&id=7251 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    CLOSED Percent Complete Details There is an issue on the SFR collect on the Paris LNS Customers have been allocated to another LNS Date Friday 04 July 2014 19 10PM Reason for closing Done Comments 1 Comment by OVH Friday 04 July 2014 19 10PM SFR have had an incident on the line in the TH2 site The incident has been resolved since 18 20 French time Customers PPP sessions

    Original URL path: http://status.ovh.es/?do=details&id=7215 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Incident Category Backend Core Status CLOSED Percent Complete Details Our collect provider has an incident on a BAS Accesses are synchronized but do not have connection Everything should get back to normal in 1H A restart of the modem may

    Original URL path: http://status.ovh.es/?do=details&id=7189 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details There has been an incident on the national FT collect gateway Sessions have all been remounted on the backup LNS Date Friday 20 June 2014 13 38PM Reason for closing Done Comments 2 Comment by OVH Friday 20 June 2014 13 34PM The gateway is now operational in Roubaix sessions will switched over to the LNS Comment by OVH Friday 20 June

    Original URL path: http://status.ovh.es/?do=details&id=7118 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    On the night of Saturday May 31st to Sunday June 1st we will switch clients sessions in collecte FT on regional LNS to a national LNS Paris or Roubaix A cut off is expected during the operation Date Sunday 01

    Original URL path: http://status.ovh.es/?do=details&id=6988 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    We are carrying out maintenance on the maintenance on the optical loop in the north of Petit Couronne We are adding 10G services Securisation will be used to protect against any outages Date Monday 19 May 2014 16 19PM Reason

    Original URL path: http://status.ovh.es/?do=details&id=6915 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    May 2014 01 40AM Reason for closing Done Comments 5 Comment by OVH Friday 16 May 2014 16 25PM The ANJ75 and GUT75 DSLAMs are no longer available We will investigate Comment by OVH Friday 16 May 2014 16 26PM The ANJ75 and GUT75 DSLAMs are available again since 15 45 CET We re working to restore the security of the optical loop Comment by OVH Friday 16 May 2014

    Original URL path: http://status.ovh.es/?do=details&id=6910 (2015-10-15)
    Open archived version from archive



  •